OpenStudy (studygeek):

what effect will increasing wave speed have on frequency, wavelength, and amplitude?

OpenStudy (jamesj):

It is really going to depend on the physics of the particular situation and how we actually measure these two features. For example, light passing through a medium, such as glass slows down as the photons interact with the atoms. This manifests with the wavelength decreasing. Sound, when it passes from air to water speeds up, because the atoms of the water are denser than air and transmit the vibrations more quickly. But the effect on frequency are quite complicated, and can both increase and decrease under different conditions. See, for example, https://ccrma.stanford.edu/~blackrse/h2o.html

OpenStudy (savvy):

since v = f.w f: frequency w:wavelength so if v increases, both can increase or just one can increase and they have no effect on amplitude...

OpenStudy (anonymous):

v=fw f<v f=1/t w=v/f v=aplitude